OpenStudy (ray_sihota):

I am having difficulty finding 3 points for 3x-2y=0. Please help.

OpenStudy (anonymous):

pick a value for x and then solve for y. the pair (x, y) give a point.

OpenStudy (anonymous):

you there?

OpenStudy (ray_sihota):

yep just trying out the equation so 0 is the value for x 3(0)-2y=0?

OpenStudy (anonymous):

so what does y have to be?

OpenStudy (ray_sihota):

y would have to be 0

OpenStudy (anonymous):

excellent! so now you have one point, (0, 0). pick another value for x

OpenStudy (ray_sihota):

lets use 1 3(1)=2y-0 x=3

OpenStudy (anonymous):

hold on there... if x = 1, then 3(1)-2y=0 => 2y=3 so what is y?

OpenStudy (ray_sihota):

3x-2y=0 3(1) =3 -2y=0 x=3 that is the part I am stuck on?

OpenStudy (ray_sihota):

so the y would be 3/2 so it will not work for graphing

OpenStudy (anonymous):

first off, we specifiy x so we already know what it is. we specifiy that x = 1, this gives us... 3x - 2y = 0 so with x = 1 => 3(1) - 2y = 0 => 3 - 2y = 0 the goal here is to find out what of y goes with x = 1. and yes, is see you have y = 3/2. Awesome!

OpenStudy (anonymous):

if you don't want fractions, use multiples of 2 for your x values.

OpenStudy (ray_sihota):

3(2)-2y=0 x=6 y =6/-2= -3

OpenStudy (anonymous):

3(2)-2y=0 => 2y=6 and y = 3

OpenStudy (ray_sihota):

Is there a limit of values that I can use for x,y ,

OpenStudy (ray_sihota):

how do you know when to use a negative or positive, is that dependent on the if the line is going straight or not

OpenStudy (anonymous):

no, just pick a value for x and you can find the corresponding values of y. in fact, you can solve the equation for y, which will directly specify the y value for any given x value: 3x-2y=0 now add 2y to both sides and get 3x = 2y + 0 = 2y => 2y = 3x (equality) now divide bith sides by 2 and get \[y=\frac{ 3 }{ 2 }x\]

OpenStudy (anonymous):

so when x = 0 => y = 3/2 (0) = 0 => (x, y) = (0, 0) when x = 2 => y = 3/2 (2) = 3 => (x, y) = (2, 3) pick another value for x... preferably a multiple of 2.

OpenStudy (ray_sihota):

if I chose -2 -3(-2)=6 6/-2= -3 x=-2, y=-3

OpenStudy (ray_sihota):

thanks man appreciate the help, this section was driving me bonkers. how do i give you a medel

OpenStudy (anonymous):

very good! sorry was talking to someone here. just click on the "best response" button, but no worries. just happy to help!

OpenStudy (ray_sihota):

I appreciate it alot

OpenStudy (anonymous):

you're welcome!!!